# Provable Benefits of Unsupervised Pre-training and Transfer Learning via Single-Index Models

### v1

@@ 1 Introduction
A prominent setting with additional data is semi-supervised learning, where one might have an abundance of unlabeled data. Unsupervised pre-training has emerged as a popular strategy in this context (Devlin (2018), Brown (2020)). The core idea behind pre-training is to train a model (on the unlabeled data) on a task that is related or, even better, a necessary precursor to the supervised task of interest. Initializing from a pre-trained model, the hope is that the model will have already learned useful features from the unlabeled data and hence solve the supervised task with reduced sample complexity. The model trained in this way can then be used to initialize a model for the supervised task in various ways e.g., by removing the final layer of the network and replacing it with an output head relevant to the labeled task, such as a classification head.
Another prominent setup with additional data is transfer learning Pan and Yang (2009), where one has access to samples from related supervised tasks. In this case, a natural idea is to initialize the model weights on the target dataset using the model trained on the upstream task.
The precise scheme for unsupervised pre-training or transfer learning can vary significantly across applications. For example, BERT models employ self-supervised representation learning by predicting “masked” tokens based on the observed ones Devlin (2018). Similar self-supervised pre-training algorithms form core components in the training of modern language models such as GPT (Radford et al. (2019), Brown (2020), Achiam et al. (2023)) and have attracted widespread attention recently. Many other forms of both pre-training and transfer leaning are applied in countless works spanning many different fields within machine learning (Wang et al. (2016) , He et al. (2017), Devlin (2018), Hagos and Kant (2019), Schneider et al. (2019)). Their popularity underscores the importance of understanding the effects of these different methods of weight initialization on supervised tasks.
@@ 2.1 Pre-training and Transfer Learning Theory
There has been significant recent progress in understanding the benefits of distinct unsupervised pre-training methods. In Lee et al. (2021), the authors provide rigorous evidence of the benefits of self-supervised pre-training (SSL). They explain the benefits of SSL via specific conditional independence relations between two sets of observed features, given the response. In a related direction, Arora et al. (2019); Tosh et al. (2021a, b) examine the benefits of contrastive pretraining, while Zhai et al. (2023) examines the effects of augmentation-based self-supervised representation learning. In Wei et al. (2021), the authors explore the benefits pre-trained language models, while Zhang and Hashimoto (2021), explores the inductive bias of masked language modeling by connecting to the statistical literature on learning graphical models. Finally, we highlight the work Azar and Nadler (2024), which exhibits provable computational benefits of semi-supervised learning under the low-degree likelihood hardness conjecture.
The paucity of high-quality labeled data has directly motivated inquiries into the properties of transfer learning across diverse application domains. The recent literature focuses on several distinct notions of transfer learning (e.g., covariate shift Heckman (1979); Huang et al. (2006); Shimodaira (2000), model shift Wang and Schneider (2015); Wang et al. (2014), target shift Maity et al. (2022), conditional shift Quiñonero-Candela et al. (2022); Storkey (2008) etc) and develops distinct rigorous methods to ensure successful knowledge transfer in these settings (see Shimodaira (2000); Wu et al. (2019); Sagawa et al. (2019); Ganin et al. (2016); Long et al. (2017) and the references therein for an incomplete list). From a learning theoretic perspective, recent works study the generalization performance as a function of the discrepancy between the source and the target domains Albuquerque et al. (2019); Ben-David et al. (2010); David et al. (2010); Hanneke and Kpotufe (2019); Tachet des Combes et al. (2020); Zhao et al. (2019).
@@ 2.2 Understanding Sample Complexity for single-index models
Single-index models have emerged as popular toy-models for understanding the sample complexity of training of neural networks. This is due to the fact that they are both high-dimensional and non-convex. From a statistical perspective there has been work on the fundamental thresholds of inference in these problems Barbier et al. (2019); Maillard et al. (2020) and its landscape geometry Sun et al. (2018); Maillard et al. (2019); Dudeja and Hsu (2018). From the perspective of sample complexity, a substantial amount of deep work in this direction focused on the sample complexity of spectral methods or related algorithms, particularly in relation to the Phase Retrieval problem, Candes et al. (2015); Barbier et al. (2019); Lu and Li (2020).
More recently there have been tight analyses of the sample complexity for online stochastic gradient descent from random initialization. In particular, it was shown in Ben Arous et al. (2021) that the sample complexity in the online setting is characterized by the Information Exponent. Since then there has been a tremendous body of work around complexity exponents, such as the Information Exponent, Leap Exponent Abbe et al. (2023), or Generative Exponent Damian et al. (2024b). In particular, these exponents have enabled studies which contrast the performance of various learning paradigms such as Correlational Statistical Query (CSQ) versus Statistical Query (SQ) bounds Damian et al. (2024b), feature learning versus kernel methods Ba et al. (2024), better choices of loss function Damian et al. (2024a), and the importance of data reuse Dandi et al. (2024); Lee et al. (2024). We note here that there has been quite a lot of recent important work on the case of multi-index models which we do not explore here, see, e.g., Abbe et al. (2023); Bietti et al. (2023); Ren and Lee (2024) for a small selection of this rich literature.
To our knowledge, most of this work has focused largely on the setting of isotropic Gaussian features (though note Zweig et al. (2024) for work on universality). However, given that pre-training only has access to the features, one requires that the features have some correlation with the underlying spike. Inspired by the recent works of Mousavi-Hosseini et al. (2023); Ba et al. (2024), we model this via a spiked covariance model.
@@ 3.1 Problem Set Up and Notation
Let the labeled data be $(y_{i},a_{i})_{i=1}^{N}$, with each $(y_{i},a_{i})$ independent and identically distributed. We have the following relationship between $a$ and $y$: $y_{i}=f(a_{i}\cdot v_{0})+\epsilon_{i}$, for some $\epsilon_{i}$ independent of $a_{i}$, with mean 0 and finite fifth moment. The parameter vector we wish to estimate is $v_{0}$ and $f$ is a known activation function. Throughout we assume that $f$ is twice differentiable almost everywhere with $f,f^{\prime},f^{\prime\prime}$ of at most polynomial growth. We would like the model we consider to capture the essence of pre-training. To perform pre-training, one should have access to additional unlabeled data. We thus assume access to some unlabeled $(a_{i}^{\prime})_{i=1}^{N^{\prime}}$ with $a_{i}^{\prime}\stackrel{{\scriptstyle D}}{{=}}a_{i}$. In order for pre-training to be useful, there is an implicit assumption that the unlabeled data contains some information in its structure that is related to the supervised task. We thus let $a_{i}\sim N(0,I_{d}+\lambda vv^{T})$, with $v\cdot v_{0}=\eta_{1}\in[0,1]$ and $v,v_{0}\in\mathbb{S}^{d-1}$ (the unit sphere in $\mathbb{R}^{d}$). Thus, the features are Gaussian with spiked covariance, where the spike vector has some correlation with the unknown parameter vector of interest $v_{0}$. In this way, our model captures the significance of pre-training by allowing the unlabeled feature data to contain information hidden in its covariance structure that is directly correlated with the solution of the supervised learning problem. The value of $\eta_{1}$ measures the strength of this correlation. We define $\eta_{2}=\sqrt{1-\eta_{1}^{2}}$.
@@
We want to understand the benefit of pre-training and so we consider two methods of initializations, random and with pre-training. For random initializations we let $X_{0}\sim\mathrm{Uniform}(\mathbb{S}^{d-1})$. To model pre-training, we use Principal Component Analysis (PCA) on the unlabeled data $(a_{i}^{\prime})_{i=1}^{N^{\prime}}$, to obtain an estimate $\hat{v}$ of the spike direction $v$. We then use this to initialize SGD for our supervised task, that is we let $X_{0}=\hat{v}$. PCA is arguably the ideal starting point for a rigorous investigation into unsupervised pre-training. From a statistical perspective, PCA is the simplest dimension-reduction algorithm; further, it’s properties are well-understood in high-dimensions Bai and Silverstein (2010). More importantly, it has been shown that more advanced representation learning algorithms, such as reconstruction autoencoders, also implement PCA in certain regimes Bourlard and Kamp (1988); Baldi and Hornik (1989); Nguyen (2021). In this light, we will restrict ourselves to PCA-based pre-training in this paper. We note that this approach is distinct from the spectral methods introduced for single-index models and, in particular, phase retrieval. There, the methods are supervised in that they use both knowledge of the label and features, whereas in unsupervised pre-training one only has access to the features.
@@ 3.2 Main Results
for all $X$ such that $(m_{1}(X),m_{2}(X))\in B_{2}(0,1):m_{1}(X)\geq m_{1}^{*},\left|m_{2}(X)\right|<m_{2}^{*}$.
We note that the ability of $\Phi$ to meet Assumption 3.2 depends entirely on the choice of activation function $f$. It is clear that population flow, initialized within the rectangle defined by the points $m^{*}$ and $(1,0)$ will recover the correct solution (see Figure 1 for an example), and hence this assumption provides a simple way to verify that a sequence of initializations is Effective for population flow. Now we state our main result regarding initializing with pre-training.
@@
for $g\sim\mathcal{N}(0,1)$. We now consider our second main result which considers recovery from random initializations.
@@
Suppose that $f$ satisfies (1) for $g\sim\mathcal{N}(0,1)$. When $\eta_{1}=1$, for spherical SGD with $N=\alpha d$ steps where $\alpha=\omega(1)$, $\alpha\delta^{2}\ll d^{1/3}$ , we have that there exists some $r>0$ such that for all sequences of initializations $X_{0}^{(d)}\in H_{r}^{(d)}$, then we have that: $m_{1}(X_{N})\to 0$ in probability as $d\to\infty$. Further we have that for all $\epsilon>0$:
@@
This result demonstrates the surprising fact that in the simple scenario where the spike is perfectly aligned with the unknown parameter vector, that with any amount of data, given appropriate stepsize, recovery is not possible from random initializations. Even more surprising, not only is recovery not possible from random initialization, but even initializing with some fixed correlation, can result in not only a failure to recover but further a loss of the initial correlation. We also have that the maximum correlation attained over the course of SGD is contained in a ball around 0 with radius slightly larger than the ball containing the initializations. Taking into account Theorem 3.3, we see that there exists problems such that pre-training can allow us to solve the problem in linear time, whereas the problem is unsolvable from random initialization in the given scaling regime, regardless of the amount of labeled data.
@@ 4 Transfer learning
The proof of this theorem follows almost exactly from Ben Arous et al. (2021), noting that their arguments still hold under slightly different initializations. Contrasting this theorem with the results of Ben Arous et al. (2021), we notice that when $v^{(d)}\cdot v_{0}^{(d)}=\eta_{d}=O(d^{-\zeta})$ for some $\zeta\in(0,1/2)$ and the information exponent is 3 or greater, we benefit from a polynomial sample reduction from $\alpha\gg d^{k-2}(\log d)^{2}$ to $\alpha\gg d^{2\zeta(k-2)}(\log d)^{2}$. Further in the case that $\zeta=0$, i.e., we initialize with a fixed correlation independent of the dimension, we see that only $\alpha=\omega(1)$ is required, and hence we can recover in nearly linear sample complexity regardless of the information exponent. This offers a substantial polynomial reduction in the event that the information exponent is large. We do note however that even in the case the of the information exponent 2, we still benefit from a complexity reduction of a factor of $(\log d)^{2}$. In the case of information exponent 1, we do not benefit from transfer learning, however, in this case we already recover with nearly linear sample complexity, hence we have no need to perform transfer learning.
@@ 7 Conclusion
Acknowledgements
A.J. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada Research Chairs programme, and the Ontario Research Fund. Cette recherche a été enterprise grâce, en partie, au soutien financier du Conseil de Recherches en Sciences Naturelles et en Génie du Canada (CRSNG), [RGPIN-2020-04597, DGECR-2020-00199], et du Programme des chaires de recherche du Canada. S.S. gratefully acknowledges support from NSF (DMS CAREER 2239234), ONR (N00014-23-1-2489) and AFOSR (FA9950-23-1-0429).
@@ Appendix A Proofs of Main Results
See the proof of proposition B.1 in Ben Arous et al. [2021] and not that similar arguments apply when the features have spiked covariance. ∎
@@
We will show that the spherical projections are negligible (arbitrarily small with probability 1 - o(1)). We can then consider the linearized paths of $(m_{1}(X_{t}),m_{2}(X_{t}))$ and $(m_{1}(\bar{X_{t}}),m_{2}(\bar{X_{t}}))$. Bounding their difference with Gronwall’s inequality Gross [1967] and Doob’s Inequality Williams [1991] to control the martingale term.
@@
with probability $1-o(1)$, by applying Doob’s inequality and noting the use of (3). Thus applying the discrete Gronwall inequality Gross [1967], we obtain:
@@
with probability $1-o(1)$. This follows from another application of Doob’s inequality onto the projection of the high dimensional martingale onto the two fixed directions of interest $e_{1},e_{2}$ keeping in mind Lemma A.1. We have thus shown that for any $\gamma>0$, with probability $1-o(1)$, that $\sup_{t\leq N}\|(m_{1}(X_{t}),m_{2}(X_{t}))-(m_{1}(\bar{X_{t}}),m_{2}(\bar{X_{t}}))\|_{2}<\gamma$ which concludes the proof.
@@
for small enough $\gamma$. The second equality follows for large $d$ using a well known result regarding the limiting correlation of $v$ and $\hat{v}$ as can be seen in Bandeira et al. [2020].
@@
Once, again, so long as $m_{1}(X_{t})\leq 2r$, we have that $\nabla\Phi(X_{i}))\cdot v_{0}>0$. By the same contradiction based argument as before, we can show that $m_{1}(X_{t})\leq 2r$ for all $t\in[0,\min(\tau_{1},N+1))$ (recalling that $m_{1}(X_{0})<r$). Hence we have the bound:
@@
The proof of this theorem follows from Ben Arous et al. [2021]. Specifically in section 2.1 of their paper they show that the single-index model covered here meets the assumptions required for there main results which apply more generally. Then noting that by replacing a random uniform initialization of order $d^{-1/2}$ with a fixed initialization of order $d^{-\zeta}$, all their arguments for Theorem 1.3 of their paper still hold, with the new sample complexity provided above (depending on $\zeta$). In the case of $\zeta=0$, the result simply follows by applying Theorem 3.2 of their paper, noting the arguments used to prove this theorem apply just as well when considering a sequence of initializations which itself is not constant but is bounded above and below by constants. ∎

### v2

@@ 1 Introduction
A prominent setting with additional data is semi-supervised learning, where one might have an abundance of unlabeled data. Unsupervised pre-training has emerged as a popular strategy in this context Devlin et al. (2019); Brown (2020). The core idea behind pre-training is to train a model (on the unlabeled data) on a task that is related or, even better, a necessary precursor to the supervised task of interest. Initializing from a pre-trained model, the hope is that the model will have already learned useful features from the unlabeled data and hence solve the supervised task with reduced sample complexity. The model trained in this way can then be used to initialize a model for the supervised task in various ways e.g., by removing the final layer of the network and replacing it with an output head relevant to the labeled task, such as a classification head.
Another prominent setup with additional data is transfer learning Pan & Yang (2009), where one has access to samples from related supervised tasks. In this case, a natural idea is to initialize the model weights on the target dataset using the model trained on the upstream task.
The precise scheme for unsupervised pre-training or transfer learning can vary significantly across applications. For example, BERT models employ self-supervised representation learning by predicting “masked” tokens based on the observed ones Devlin et al. (2019). Similar self-supervised pre-training algorithms form core components in the training of modern language models such as GPT Radford et al. (2019); Brown (2020); Achiam et al. (2023) and have attracted widespread attention recently. Many other forms of both pre-training and transfer leaning are applied in countless works spanning many different fields within machine learning Wang et al. (2016); He et al. (2017); Devlin et al. (2019); Schneider et al. (2019). Their popularity underscores the importance of understanding the effects of these different methods of weight initialization on supervised tasks.
@@ 2.1 Pre-training and Transfer Learning Theory
There has been significant recent progress in understanding the benefits of distinct unsupervised pre-training methods. In Lee et al. (2021), the authors provide rigorous evidence of the benefits of self-supervised pre-training (SSL). They explain the benefits of SSL via specific conditional independence relations between two sets of observed features, given the response. In a related direction, Arora et al. (2019); Tosh et al. (2021a, b) examine the benefits of contrastive pretraining, while Zhai et al. (2023) examines the effects of augmentation-based self-supervised representation learning. In Wei et al. (2021), the authors explore the benefits pre-trained language models, while Zhang & Hashimoto (2021), explores the inductive bias of masked language modeling by connecting to the statistical literature on learning graphical models. Finally, we highlight the work Azar & Nadler (2025), which exhibits provable computational benefits of semi-supervised learning under the low-degree likelihood hardness conjecture.
The paucity of high-quality labeled data has directly motivated inquiries into the properties of transfer learning across diverse application domains. The recent literature focuses on several distinct notions of transfer learning (e.g., covariate shift Heckman (1979); Huang et al. (2006); Shimodaira (2000), model shift Wang & Schneider (2015); Wang et al. (2014), target shift Maity et al. (2022), conditional shift Quiñonero-Candela et al. (2022); Storkey (2008) etc) and develops distinct rigorous methods to ensure successful knowledge transfer in these settings (see Shimodaira (2000); Ganin et al. (2016); Long et al. (2017); Wu et al. (2019); Sagawa et al. (2019); Gerace et al. (2022) and the references therein for an incomplete list). From a learning theoretic perspective, recent works study the generalization performance as a function of the discrepancy between the source and the target domains Albuquerque et al. (2019); Ben-David et al. (2010); David et al. (2010); Hanneke & Kpotufe (2019); Tachet des Combes et al. (2020); Zhao et al. (2019).
@@ 2.2 Understanding Sample Complexity for single-index models
Single-index models have emerged as popular toy-models for understanding the sample complexity of training of neural networks. This is due to the fact that they are both high-dimensional and non-convex. From a statistical perspective there has been work on the fundamental thresholds of inference in these problems Barbier et al. (2019); Maillard et al. (2020b) and its landscape geometry Sun et al. (2018); Maillard et al. (2020a); Dudeja & Hsu (2018). From the perspective of sample complexity, a substantial amount of deep work in this direction focused on the sample complexity of spectral methods or related algorithms, particularly in relation to the Phase Retrieval problem, Candes et al. (2015); Barbier et al. (2019); Lu & Li (2020).
More recently there have been tight analyses of the sample complexity for online stochastic gradient descent from random initialization. In particular, it was shown in Ben Arous et al. (2021) that the sample complexity in the online setting is characterized by the Information Exponent. Since then there has been a tremendous body of work around complexity exponents, such as the Information Exponent, Leap Exponent Abbe et al. (2023), or Generative Exponent Damian et al. (2024b). In particular, these exponents have enabled studies which contrast the performance of various learning paradigms such as Correlational Statistical Query (CSQ) versus Statistical Query (SQ) bounds Damian et al. (2024b), feature learning versus kernel methods Ba et al. (2024), better choices of loss function Damian et al. (2024a), and the importance of data reuse Dandi et al. (2024); Lee et al. (2024); Arnaboldi et al. (2024). We note here that there has been quite a lot of recent important work on the case of multi-index models which we do not explore here, see, e.g., Abbe et al. (2023); Bietti et al. (2023); Ren & Lee (2024) for a small selection of this rich literature.
To our knowledge, most of this work has focused largely on the setting of isotropic Gaussian features (though note Zweig et al. (2024); Pesce et al. (2023) for work on universality). However, given that pre-training only has access to the features, one requires that the features have some correlation with the underlying spike. Inspired by the recent works of Mousavi-Hosseini et al. (2023); Ba et al. (2024), we model this via a spiked covariance model. There are some additional recent works citedandi2024random, cui2024asymptotics, ba2022high which analyze the high dimensional asymptotics of neural network training after a single gradient step and also consider spiked and isotropic index models.
@@ 3.1 Problem Set Up and Notation
Let the labeled data be $(y_{i},a_{i})_{i=1}^{N}$, with each $(y_{i},a_{i})$ independent and identically distributed. We have the following relationship between $a$ and $y$: $y_{i}=f(a_{i}\cdot v_{0})+\epsilon_{i}$, for some $\epsilon_{i}$ independent of $a_{i}$, with mean 0 and finite fifth moment. The parameter vector we wish to estimate is $v_{0}$ and $f$ is a known activation function. Throughout we assume that $f$ is twice differentiable almost everywhere with $f,f^{\prime},f^{\prime\prime}$ of at most polynomial growth. We would like the model we consider to capture the essence of pre-training. To perform pre-training, one should have access to additional unlabeled data. We thus assume access to some unlabeled $(a_{i}^{\prime})_{i=1}^{N^{\prime}}$ with $a_{i}^{\prime}\stackrel{{\scriptstyle D}}{{=}}a_{i}$. In order for pre-training to be useful, there is an implicit assumption that the unlabeled data contains some information in its structure that is related to the supervised task. We thus let $a_{i}\sim N(0,I_{d}+\lambda vv^{T})$, with $v\cdot v_{0}=\eta_{1}\in[0,1],\lambda>0$ ($\eta_{1},\lambda$ are dimension independent) and $v,v_{0}\in\mathbb{S}^{d-1}$ (the unit sphere in $\mathbb{R}^{d}$). Thus, the features are Gaussian with spiked covariance, where the spike vector has some correlation with the unknown parameter vector of interest $v_{0}$. In this way, our model captures the significance of pre-training by allowing the unlabeled feature data to contain information hidden in its covariance structure that is directly correlated with the solution of the supervised learning problem. The value of $\eta_{1}$ measures the strength of this correlation. We define $\eta_{2}=\sqrt{1-\eta_{1}^{2}}$.
@@
We want to understand the benefit of pre-training and so we consider two methods of initializations, random and with pre-training. For random initializations we let $X_{0}\sim\mathrm{Uniform}(\mathbb{S}^{d-1})$. To model pre-training, we use Principal Component Analysis (PCA) on the unlabeled data $(a_{i}^{\prime})_{i=1}^{N^{\prime}}$, to obtain an estimate $\hat{v}$ of the spike direction $v$. We then use this to initialize SGD for our supervised task, that is we let $X_{0}=\hat{v}$. PCA is arguably the ideal starting point for a rigorous investigation into unsupervised pre-training. From a statistical perspective, PCA is the simplest dimension-reduction algorithm; further, it’s properties are well-understood in high-dimensions Bai & Silverstein (2010). More importantly, it has been shown that more advanced representation learning algorithms, such as reconstruction autoencoders, also implement PCA in certain regimes Bourlard & Kamp (1988); Baldi & Hornik (1989); Nguyen (2021). In this light, we will restrict ourselves to PCA-based pre-training in this paper. We note that this approach is distinct from the spectral methods introduced for single-index models and, in particular, phase retrieval. There, the methods are supervised in that they use both knowledge of the label and features, whereas in unsupervised pre-training one only has access to the features.
@@ 3.2 Main Results
for all $X$ such that $(m_{1}(X),m_{2}(X))\in B_{2}(0,1):m_{1}(X)\geq m_{1}^{*},\left|m_{2}(X)\right|<m_{2}^{*}$. Where $B_{2}(0,1)$ is the 2-d unit ball.
We note that the ability of $\Phi$ to meet Assumption 3.2 depends entirely on the choice of activation function $f$. It is clear that population flow, initialized within the rectangle defined by the points $m^{*}$ and $(1,0)$ will recover the correct solution (see Figure 1 for an example), hence (emphasizing that $m^{*}$ is independent of the dimension) this assumption provides a simple way to verify that a sequence of initializations is Effective for population flow. Now we state our main result regarding initializing with pre-training.
@@
for $g\sim\mathcal{N}(0,1)$. Recalling the notion of information exponent from Ben Arous et al. (2021) which is restated in the following section, (1) is equivalent to requiring the information exponent 3 or greater for $f$ and the information exponent less than or equal to 2 for $f^{2}$. We now consider our second main result which considers recovery from random initializations.
@@
Suppose that $f$ satisfies (1) for $g\sim\mathcal{N}(0,1)$. When $\eta_{1}=1$, for spherical SGD with $N=\alpha d$ steps where $\alpha=\omega(1)$, $\alpha\delta^{2}\ll d^{1/3}$ , we have that there exists some dimension independent value $r>0$ such that for all sequences of initializations $X_{0}^{(d)}\in H_{r}^{(d)}$, then we have that: $m_{1}(X_{N})\to 0$ in probability as $d\to\infty$. Further we have that for all $\epsilon>0$:
@@
This result demonstrates the surprising fact that in the simple scenario where the spike is perfectly aligned with the unknown parameter vector, recovery is not possible from random initializations with any amount of data, given appropriate stepsize. Even more surprising, not only is recovery not possible from random initialization, but even initializing with some fixed correlation, can result in not only a failure to recover but further a loss of the initial correlation. We also have that the maximum correlation attained over the course of SGD is contained in a ball around 0 with radius slightly larger than the ball containing the initializations. Taking into account Theorem 3.3, we see that there exists problems such that pre-training can allow us to solve the problem in linear time, whereas the problem is unsolvable from random initialization in the given scaling regime, regardless of the amount of labeled data.
@@ 4 Transfer learning
The proof of this theorem follows almost exactly from Ben Arous et al. (2021), noting that their arguments still hold under slightly different initializations. Contrasting this theorem with the results of Ben Arous et al. (2021), we notice that when $v^{(d)}\cdot v_{0}^{(d)}=\eta_{d}=O(d^{-\zeta})$ for some $\zeta\in(0,1/2)$ and the information exponent is 3 or greater, we benefit from a polynomial sample reduction from $\alpha\gg d^{k-2}(\log d)^{2}$ to $\alpha\gg d^{2\zeta(k-2)}(\log d)^{2}$. Further in the case that $\zeta=0$, i.e., we initialize with a fixed correlation independent of the dimension, we see that only $\alpha=\omega(1)$ is required, and hence we can recover in nearly linear sample complexity regardless of the information exponent. This offers a substantial polynomial reduction in the event that the information exponent is large. We do note however that even in the case the of the information exponent 2, we still benefit from a complexity reduction of a factor of $(\log d)^{2}$. In the case of information exponent 1, we do not benefit from transfer learning, however, in this case we already recover with nearly linear sample complexity, hence we have no need to perform transfer learning. We also note here that the reduction in sample complexity is the same as observed in Theorem 7 of Mousavi-Hosseini et al. (2023).
@@ 7 Conclusion
Acknowledgments
A.J. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada Research Chairs programme, and the Ontario Research Fund. Cette recherche a été enterprise grâce, en partie, au soutien financier du Conseil de Recherches en Sciences Naturelles et en Génie du Canada (CRSNG), [RGPIN-2020-04597, DGECR-2020-00199], et du Programme des chaires de recherche du Canada. S.S. gratefully acknowledges support from NSF (DMS CAREER 2239234), ONR (N00014-23-1-2489) and AFOSR (FA9950-23-1-0429). T.J acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).
Impact Statement
This paper focuses on advancing the theoretical foundations of Machine Learning, and we do not anticipate significant societal consequences from our work.
@@ Appendix A Proofs of Main Results
See the proof of proposition B.1 in Ben Arous et al. (2021) and not that similar arguments apply when the features have spiked covariance. ∎
@@
We will show that the spherical projections are negligible (arbitrarily small with probability 1 - o(1)). We can then consider the linearized paths of $(m_{1}(X_{t}),m_{2}(X_{t}))$ and $(m_{1}(\bar{X_{t}}),m_{2}(\bar{X_{t}}))$. Bounding their difference with Gronwall’s inequality Gross (1967) and Doob’s Inequality Williams (1991) to control the martingale term.
@@
with probability $1-o(1)$, by applying Doob’s inequality and noting the use of (3). Thus applying the discrete Gronwall inequality Gross (1967), we obtain:
@@
with probability $1-o(1)$. This follows from another application of Doob’s inequality onto the projection of the high dimensional martingale onto the two fixed directions of interest $e_{1},e_{2}$ keeping in mind Lemma A.1. We have thus shown that for any $\gamma>0$, with probability $1-o(1)$, that $\sup_{t\leq N}\|(m_{1}(X_{t}),m_{2}(X_{t}))-(m_{1}(\bar{X_{t}}),m_{2}(\bar{X_{t}}))\|_{2}<\gamma$ which concludes the proof. ∎
@@
for small enough $\gamma$. The second equality follows for large $d$ using a well known result regarding the limiting correlation of $v$ and $\hat{v}$ as can be seen in Bandeira et al. (2020).
@@
Once again, so long as $m_{1}(X_{t})\leq 2r$, we have that $\nabla\Phi(X_{i}))\cdot v_{0}>0$. By the same contradiction based argument as before, we can show that $m_{1}(X_{t})\leq 2r$ for all $t\in[0,\min(\tau_{1},N+1))$ (recalling that $m_{1}(X_{0})<r$). Hence we have the bound:
@@
The proof of this theorem follows from Ben Arous et al. (2021). Specifically in section 2.1 of their paper they show that the single-index model covered here meets the assumptions required for there main results which apply more generally. Then noting that by replacing a random uniform initialization of order $d^{-1/2}$ with a fixed initialization of order $d^{-\zeta}$, all their arguments for Theorem 1.3 of their paper still hold, with the new sample complexity provided above (depending on $\zeta$). In the case of $\zeta=0$, the result simply follows by applying Theorem 3.2 of their paper, noting the arguments used to prove this theorem apply just as well when considering a sequence of initializations which itself is not constant but is bounded above and below by constants. ∎